# Evidence of 1:1 slope between rocky Super-Earths and their host stars

### v1

@@ 2.4 Statistical method
This linear best-fit yields an empirical relationship that may be interpreted in the context of formation scenarios. However, a different method to determine whether the planet-star data supports different formation scenarios is to use a Hierarchical Bayesian Modeling (HBM) that ties all the data at a population level. After running some tests to obtain hyperparameters that could describe a systematic shift from or an intrinsic scatter around stellar composition due to formation reprocessing, we deemed the sample is too small and the errors are too large at the moment. This is worth revisiting once enough data is acquired (for example after PLATO launches PLATO).
In addition to a sound statistical treatment, we need to pay careful attention to planets that straddle the volatile-rocky boundary (e.g. rocky threshold radius–RTR or largest size for an MgSiO3+MgO planet) as well as the pure iron line. These exoplanets have part of their probably density functions in M-R space within the rocky region and part of it outside. One way to model them is to restrict the composition solutions only to the rocky subset (employed by Plotnykov2020 and Adibekyan2021), but this artificially increases the certainty in their composition leading to small error bars that would anchor any inferred relationship slope. Without the knowledge on the absence of an envelope, this method may yield misleading results. Another way is to allow planets to have negative CMFs to represent the M-R solutions that call for volatiles (Brinkman2024). This is done by extrapolating fits to that are obtained within the rocky region, to negative values. This, of course, is non-physical and highly dependent on the extrapolation scheme.
@@
3 Results
3.1 Link between planets and stars
To compare compositions of exoplanets to their host stars (planet-star) we apply our most sophisticated planetary interior model with full mineralogy consideration and our chemical model that translates stars refractory abundances to planet equivalent parameters (see Sec. 2.1). We search for correlations in various different parameters and focus on the refractory ratios (Fe/Si, Fe/Mg, etc.), Fe-MF, age of the system and normalized density with the goal of identifying any links or compositional trends. Note that our fiducial dataset assumes non-local thermal equilibrium (non-LTE) correction for stellar abundances and hence Fe-MFst.
We present the resulting planet-star compositions in Table 1 and include for comparison planetary CMF results that are not physically self-consistent (Plotnykov2020). Given the updated radius of 55 Cnc e, this planet is no longer considered rocky as it is above the rocky-threshold-radius (RTR; ) and may have a significant amount of volatiles, making it impossible to constrain the rocky interior. Additionally, we identify two more planets that seem to have volatiles (more on this in paperI), WASP-47 e and Kepler-36 b, given the fact that they straddle the RTR boundary. As explained, we weight the planets contributions to the fits according to how much their M-R data intersects the rocky versus volatile/over-dense regions.
3.1.1 Relationship in chemical ratios
We first compare the updated planetary data in a population sense and confirm our observation in Plotnykov2020 that exoplanets span a wider refractory ratio space (Fe/Mg, Fe/Si by weight) compared to their host stars (see Fig. 1). However, unlike the Plotnykov2020 sample, this updated data, derived from a homogeneous and physically self-consistent stellar treatment consists of a higher concentration of exoplanets around an Earth-like composition, in better agreement with a stellar composition (see Fig. 12).
@@ 3.1.1 Relationship in chemical ratios
3.1.2 Relationship in iron-mass fraction
The comparison in Fe-MF space is the most robust and direct metric for testing the composition link between planets and their host stars. With our weighted dataset, we find an intriguing possible 1:1 slope in Fe-MF space, given a slope of and an intercept of , see Figure 3. Furthermore, we do not find strong evidence for any sub-populations, either in iron-rich or iron-poor exoplanets across the sampled host stars in agreement with Brinkman2024 and Adibekyan2024. Notably, treating each exoplanet with equal weights shifts the slope away from a 1:1 relationship (see Sec. 3.2). Thus, we find the choice of statistical methods to be highly important and without proper outlier treatment, the constraints on this relationship may be biased towards steeper/shallower best-fit lines.
Altogether for the population, our data and analysis hints at similar refractory composition between stars and their planets with a 1:1 relationship, but given the large errors in planetary masses and radii, this result is still inconclusive as a zero slope () is also allowed within a confidence interval. This stands in contrast to the steep slope obtained by Adibekyan2021, Adibekyan2024 and Brinkman2024 using ODR, with slopes that have . We discuss the multiple reasons for these discrepancies between the studies in section 3.3.
3.1.3 Relationship in normalized density
@@ 3.2 Best-fit robustness
3.3 Methodology comparisons
Using a comprehensive interior structure and sound statistical modeling on a homogeneously and physically self-consistent exoplanet-stellar data, we obtain a planet-star compositional link that is close to 1:1. These results stand in contrast to the steep relationships obtained by Adibekyan2021, Adibekyan2024 and Brinkman2024. The reason for these discrepant results is due to differences in: 1) retrieved stellar abundances, 2) exoplanet M-R data arising from differences in observed stellar M-R, 3) transit data selection (more so than radial velocity data) and/or 4) the statistical and outlier consideration treatment.
@@ 3.3 Methodology comparisons
our stellar characterization model, that ensures self-consistency across stellar parameters (e.g. ) with thorough inspection of each absorption line feature ([Fe/H], [Mg/H], [Si/H], [Ca/H], [Al/H], [Ni/H], [O/H]), yields Mg and Si abundances that are on average higher compared to Adibekyan2024 and Brinkman2024 (see Fig. 11). These larger Mg and Si abundances translate to lower Fe-MFst values. These differences primarily stem from our self-consistent methodology and stellar abundance retrieval pipeline, which result in different elemental abundances ([Fe/H], [Mg/H], [Si/H]) as well as stellar metallicity estimates. Moreover, we account for several sources of uncertainty, including a non-LTE treatment leading to larger errors in Fe-MFst, which we deem more reliable. This means that for most stars, our results are consistent within 1 with Adibekyan2021/Adibekyan2024 and Brinkman2024 stellar abundances, but not the other way around as the other two datasets reported lower, but perhaps underestimated, errors on abundance values ([Fe/H], [Mg/H], [Si/H]). We illustrate these relationships between Fe-MFst and stellar abundances in Figure 5, with a comparison to the GALAH dataset. We observe a clear correlation between Fe-MFst and [Fe/H] for lower metallicity stars in the GALAH dataset. Thus, it is vital to understand whether the composition of rocky exoplanets around these stars is similar to or enhanced in comparison to the host star composition. However, inside the limited metallicity range of our host stars (-0.2¡[Fe/H]¡0.5), we find that none of the abundances are strongly correlated with the equivalent Fe-MFst composition as well as Mg/Si refractory ratio, opposite to Adibekyan2024 and Brinkman2024 data that have a negative slope between Fe-MFst vs Mg/Si. Furthermore, we find that our host stars have a wider range of Mg/Si ratios in comparison to others.This should not be overlooked in the case of stars, as changing Mg/Si ratio has a moderate effect on the total Fe-MFst budget, which is unlike the case for exoplanets where the Fe-MFpl is set by the M-R data and their large uncertainties (Plotnykov2024).
@@
changes in exoplanet data are mainly driven by the updates in stellar radius and transit depth, that feed into the planetary radius inference. For transit depth, we use the most precise measurements and make sure that each transit comes from the same instruments, similar to Brinkman2024. Meanwhile, Adibekyan2021 and Adibekyan2024 do not use isochrone analysis to update stellar M-R and instead adopt the most recently reported exoplanet M-R measurements taken by the same group. The different planetary radii used by Adibekyan2021 () translate to an average Fe-MF compared to our sample, but this difference is lower in the revised Adibekyan2024 dataset ( and Fe-MF). Moreover and similar to Adibekyan2024 and Brinkman2024, with revised transit radii, we find a reduction of the number ultra dense exoplanets, where 3 out of 4 exoplanets that had previously been identified as super-Mercuries with Fe-MF (Plotnykov2020; Adibekyan2021) are no longer as iron-enriched, with K2-106 b changing the most (see section 4.1). In general though, our exoplanet and stellar M-R data is mostly consistent with other datasets within , except for 55-Cnc e and K2-106 b (see Fig. 11, 12).
@@
a selection of exoplanets that are on the verge of volatile-to-rocky boundary have a profound impact on setting the composition relationship and to solve this problem we have weighted each exoplanet according (, see Sec. 2.4). As mentioned other studies instead employ non-statistical methods to address this issue: Adibekyan2021 considers the RTR planets to be rocky artificially increasing their precision, while Brinkman2024 allows for negative CMFs. Adibekyan2024 instead uses a more comprehensive, but time-consuming way to model these straddling planets based on an interior structure model that allows for throughout the planet (Luo2024). However, without atmospheric retrievals there is no guarantee that a water envelope is present or how much water has survived in the rocky interior (mantle or core), given that these exoplanets have high incident fluxes (Fulton2017). By assuming there is water in the interior of compact planets, the iron component increases to offset the added mass. Thus, assuming a rocky composition yields Fe-MF values that are on average 20 points lower, and represent a lower bound for iron content. For consistency, we recalculate the Fe-MF for the different datasets (Adibekyan2021, Brinkman2024, and Adibekyan2024) to properly compare methodologies.
@@
in this study we use a Bayesian approach to fit a best-line between exoplanet and stellar data, instead of ODR used by both Adibekyan2021 and Brinkman2024, that allows for asymmetrical errors in the data and better accounting of outliers, which are highly prevalent in these datasets. For the Fe-MF comparison, Adibekyan2024 applied different statistical methods including a Bayesian model (), but it did not account for asymmetrical errors nor is the dataset weighted to account for outliers. Brinkman2024 also applied OLS approach and obtained close to 1:1 relationship (), which in conjunction with their steep relation from ODR they interpreted as there not being enough evidence of a robust correlation. With a Bayesian approach, instead, we can inspect the posterior and identify any underlying discrepancies in the fit, thus quantifying the confidence in the relation as well as identify outliers, especially in light of a small data set. Therefore, we apply our nominal weighted fitting model (see Sec. 2.4) to the Adibekyan2024 and Brinkman2024 datasets for its added statistical advantage. We find that Brinkman2024 dataset has a weaker slope to the one they report, which is closer to 1:1 , while the Adibekyan2024 dataset has a slope that is slightly shallower than their reported slope (see Fig. 6). This result indicates that both our and Brinkman2024 datasets are fundamentally different compared to that of Adibekyan2024, which has a steeper slope. Furthermore, this is not attributed to the larger sample size of Adibekyan2024 (N=32) given the fact that when we limit our re-analysis of their data to the same exoplanets that are in our sample, we find the slope is still steep (). Thus, it is most likely due to the self-consistent stellar and planetary analysis done in this work.
3.4 Data needed to unequivocally test formation scenarios
@@
We use the GALAH dataset (GALAH) to generate a population of stars and assume that each star hosts an exoplanet with some mass and corresponding radius such that the refractory ratios are exactly the same as the host star. Subsequently, we apply the current observational error scatter to this mock population (, for stars and exoplanets), and fit this Fe-MF star-planet mock data to a line to recover the slope and intercept. We repeat this process 1000 times to find the variance in these parameters. We find that the generated mock populations are fairly similar to our self-consistent exoplanet population. Our results for simulations with 21 planets in the sample (N) are shown in Figure 7. The most likely outcome is a population with no super-Mercuries – comprising half of the outcomes – although 1 or even 2 are still possible, accounting for 35 and 12 of the outcomes, respectively. We find instead that having or RTR-like planets is equally probable and irregardless of the number of outliers, we always recover a slope that is consistent with 1 and 0. Thus, the likelihood of randomly observing a population with 1 super-Mercury and 2-3 RTR-like exoplanets out of 21, which is what we have in our sample, is . The results from this mock population indicate that the larger range observed in the composition of exoplanets compared to that of stars may be explained, at the level, by the scatter produced due to observation uncertainty. It also means, however, it is entirely possible that the spread in planetary composition is real and reflects formation pathways away from a primordial origin. A larger sample and especially the prevalence or absence of super-Mercuries is the path to determine whether the spread in composition is real or an artifact of errors.
Next we examine the sample size needed to determine unequivocally a 1:1 relation given the current level of mass and radius certainty. Unsurprisingly but unfortunately, we find that a sample size of N, like the one we have, is unlikely to recover the true relationship () given the large uncertainty in the data – have m=0 within confidence (see Fig. 8). However, we find it is possible to robustly recover the correct best-fit line back, at 95 confidence when there are 150 planets. For a higher confidence level (3 ), we find that for N=150 there are around a quarter of runs with in the posterior. This means that to confidently constrain the composition using only Fe-MF of planets and stars, a sample that of at least 7 times the current size is needed. The good news is that PLATO is expected to deliver about 1000 planets with radii of after the first four mission years (PLATO).
@@
4 Discussion
4.1 Super-Mercuries
We have shown the need to carefully select planetary masses and radius and in particular guarantee consistency in stellar parameters to ensure consistency in planetary M-R data. One of the shortcomings of using inconsistent data was the apparent existence of a group of super-Mercuries Plotnykov2020; Adibekyan2021; Schulze2021) with implications on the existence of our own Mercury. In our solar system, Mercury holds a unique position as the innermost planet with an unusually large amount of iron, with Fe-MF Hauck2013 or 2-times enriched with respect to solar, making it an outlier compared to Earth, Venus and Mars which have similar refractory ratios to that of the Sun (Anders1989; Grevesse2007; Lodders2009; Asplund2009). There have been different proposed formation mechanisms (Ebel2011; Hubbard2014; Cuello2016; Kruss2020; Aguichine2020; Johansen2022) to explain Mercury’s composition with giant impacts being the leading theory (Benz2007) albeit not without problems (Chau2018; Clement2021; Franco2022; Scora2024). It remains to be seen how unique its formation is. Thus, super-Mercuries are prime targets for understanding planetary formation and evolution theories.
In our updated dataset, the exoplanet M-R are revised such that we observe fewer super-Mercuries–only Kepler-107 c remains–and there is no evidence for a sub-population as seen by Adibekyan2021, in agreement with Brinkman2024. This suggest the prevalence of fewer super-Mercuries in general, which we attribute to our homogeneous modeling and updates to the transit and RV observations. Nevertheless, we still find a few exoplanets–K2-38 b, K2-229 b and K2-291 b–that have relative iron-enrichment compared to their host star (Fe-MF). Some of these were previously thought super-Mercuries and would be important to follow up, given that there is no current formation theory to explain even this smaller iron-enrichment for massive rocky planets (Scora2020; Scora2024).
4.2 Multi-Planet systems

### v2

@@ II.4 Statistical method
This linear best-fit yields an empirical relationship that may be interpreted in the context of formation scenarios. However, a different method to determine whether the planet-star data supports different formation scenarios is to use a Hierarchical Bayesian Modeling (HBM) that ties all the data at a population level. After running some tests to obtain hyperparameters that could describe a systematic shift from or an intrinsic scatter around stellar composition due to formation reprocessing, we deemed the sample is too small and the errors are too large at the moment. This is worth revisiting once enough data is acquired (for example after PLATO launches Rauer et al. (2024)).
In addition to a sound statistical treatment, we need to pay careful attention to planets that straddle the volatile-rocky boundary (e.g. rocky threshold radius–RTR or largest size for an MgSiO3+MgO planet) as well as the pure iron line. These exoplanets have part of their probably density functions in M-R space within the rocky region and part of it outside. One way to model them is to restrict the composition solutions only to the rocky subset (employed by Plotnykov and Valencia (2020) and Adibekyan et al. (2021)), but this artificially increases the certainty in their composition leading to small error bars that would anchor any inferred relationship slope. Without the knowledge on the absence of an envelope, this method may yield misleading results. Another way is to allow planets to have negative CMFs to represent the M-R solutions that call for volatiles (Brinkman et al., 2024). This is done by extrapolating fits to that are obtained within the rocky region, to negative values. This, of course, is non-physical and highly dependent on the extrapolation scheme.
@@
III Results
III.1 Link between planets and stars
To compare compositions of exoplanets to their host stars (planet-star) we apply our most sophisticated planetary interior model with full mineralogy consideration and our chemical model that translates stars refractory abundances to planet equivalent parameters (see Sec. II.1). We search for correlations in various different parameters and focus on the refractory ratios (Fe/Si, Fe/Mg, etc.), Fe-MF, age of the system and normalized density with the goal of identifying any links or compositional trends. Note that our fiducial dataset assumes non-local thermal equilibrium (non-LTE) correction for stellar abundances and hence Fe-MFst.
We present the resulting planet-star compositions in Table 1, 2 and include for comparison planetary CMF results that are not physically self-consistent (see Fig. 13). Given the updated radius of 55 Cnc e, this planet is no longer considered rocky as it is above the rocky-threshold-radius (RTR; ) and may have a significant amount of volatiles, making it impossible to constrain the rocky interior. Additionally, we identify two more planets that seem to have volatiles (more on this in paper I), WASP-47 e and Kepler-36 b, given the fact that they straddle the RTR boundary. As explained, we weight the planets contributions to the fits according to how much their M-R data intersects the rocky versus volatile/over-dense regions.
III.1.1 Relationship in chemical ratios
We first compare the updated planetary data in a population sense and confirm our observation in Plotnykov and Valencia (2020) that exoplanets span a wider refractory ratio space (Fe/Mg, Fe/Si by weight) compared to their host stars (see Fig. 1). However, unlike the Plotnykov and Valencia (2020) sample, this updated data, derived from a homogeneous and physically self-consistent stellar treatment consists of a higher concentration of exoplanets around an Earth-like composition, in better agreement with a stellar composition (see Fig. 12).
@@ III.1.1 Relationship in chemical ratios
III.1.2 Relationship in iron-mass fraction
The comparison in Fe-MF space is the most robust and direct metric for testing the composition link between planets and their host stars. With our weighted dataset, we find an intriguing possible 1:1 slope in Fe-MF space, given a slope of and an intercept of , see Figure 3. Furthermore, we do not find strong evidence for any sub-populations, either in iron-rich or iron-poor exoplanets across the sampled host stars in agreement with B24 and A24. Notably, treating each exoplanet with equal weights shifts the slope away from a 1:1 relationship (see Sec. III.2). Thus, we find the choice of statistical methods to be highly important and without proper outlier treatment, the constraints on this relationship may be biased towards steeper/shallower best-fit lines.
Altogether for the population, our data and analysis hints at similar refractory composition between stars and their planets with a 1:1 relationship, but given the large errors in planetary masses and radii, this result is still inconclusive as a zero slope () is also allowed within a confidence interval. This stands in contrast to the steep slope obtained by A21, A24 and B24 using ODR, with slopes that have . We discuss the multiple reasons for these discrepancies between the studies in section III.3.
III.1.3 Relationship in normalized density
@@ III.2 Best-fit robustness
III.3 Methodology comparisons
Using a comprehensive interior structure and sound statistical modeling on a homogeneously and physically self-consistent exoplanet-stellar data, we obtain a planet-star compositional link that is close to 1:1. These results stand in contrast to the steep relationships obtained by A21, A24 and B24. The reason for these discrepant results is due to differences in: 1) retrieved stellar abundances, 2) exoplanet M-R data arising from differences in observed stellar M-R, 3) transit data selection (more so than radial velocity data) and/or 4) the statistical and outlier consideration treatment.
@@ III.3 Methodology comparisons
our stellar characterization model, that ensures self-consistency across stellar parameters (e.g. ) with thorough inspection of each absorption line feature ([Fe/H], [Mg/H], [Si/H], [Ca/H], [Al/H], [Ni/H], [O/H]), yields Mg and Si abundances that are on average higher compared to A24 and B24 (see Fig. 11). These larger Mg and Si abundances translate to lower Fe-MFst values. These differences primarily stem from our self-consistent methodology and stellar abundance retrieval pipeline, which result in different elemental abundances ([Fe/H], [Mg/H], [Si/H]) as well as stellar metallicity estimates. Moreover, we account for several sources of uncertainty, including a non-LTE treatment leading to larger errors in Fe-MFst, which we deem more reliable. This means that for most stars, our results are consistent within 1 with A21/A24 and B24 stellar abundances, but not the other way around as the other two datasets reported lower, but perhaps underestimated, errors on abundance values ([Fe/H], [Mg/H], [Si/H]). We illustrate these relationships between Fe-MFst and stellar abundances in Figure 5, with a comparison to the GALAH dataset. We observe a clear correlation between Fe-MFst and [Fe/H] for lower metallicity stars in the GALAH dataset. Thus, it is vital to understand whether the composition of rocky exoplanets around these stars is similar to or enhanced in comparison to the host star composition. However, inside the limited metallicity range of our host stars (-0.2¡[Fe/H]¡0.5), we find that none of the abundances are strongly correlated with the equivalent Fe-MFst composition as well as Mg/Si refractory ratio, opposite to A24 and B24 data that have a negative slope between Fe-MFst vs Mg/Si. Furthermore, we find that our host stars have a wider range of Mg/Si ratios in comparison to others.This should not be overlooked in the case of stars, as changing Mg/Si ratio has a moderate effect on the total Fe-MFst budget, which is unlike the case for exoplanets where the Fe-MFpl is set by the M-R data and their large uncertainties (Plotnykov and Valencia, 2024).
@@
changes in exoplanet data are mainly driven by the updates in stellar radius and transit depth, that feed into the planetary radius inference. For transit depth, we use the most precise measurements and make sure that each transit comes from the same instruments, similar to B24. Meanwhile, A21 and A24 do not use isochrone analysis to update stellar M-R and instead adopt the most recently reported exoplanet M-R measurements taken by the same group. The different planetary radii used by A21 () translate to an average Fe-MF compared to our sample, but this difference is lower in the revised A24 dataset ( and Fe-MF). Moreover and similar to A24 and B24, with revised transit radii, we find a reduction of the number ultra dense exoplanets, where 3 out of 4 exoplanets that had previously been identified as super-Mercuries with Fe-MF (Plotnykov and Valencia, 2020; Adibekyan et al., 2021) are no longer as iron-enriched, with K2-106 b changing the most (see section IV.1). In general though, our exoplanet and stellar M-R data is mostly consistent with other datasets within , except for 55-Cnc e and K2-106 b (see Fig. 11, 12).
@@
a selection of exoplanets that are on the verge of volatile-to-rocky boundary have a profound impact on setting the composition relationship and to solve this problem we have weighted each exoplanet according (, see Sec. II.4). As mentioned other studies instead employ non-statistical methods to address this issue: A21 considers the RTR planets to be rocky artificially increasing their precision, while B24 allows for negative CMFs. A24 instead uses a more comprehensive, but time-consuming way to model these straddling planets based on an interior structure model that allows for throughout the planet (Luo et al., 2024). However, without atmospheric retrievals there is no guarantee that a water envelope is present or how much water has survived in the rocky interior (mantle or core), given that these exoplanets have high incident fluxes (Fulton et al., 2017). By assuming there is water in the interior of compact planets, the iron component increases to offset the added mass. Thus, assuming a rocky composition yields Fe-MF values that are on average 20 points lower, and represent a lower bound for iron content. For consistency, we recalculate the Fe-MF for the different datasets (A21, B24, and A24) to properly compare methodologies.
@@
in this study we use a Bayesian approach to fit a best-line between exoplanet and stellar data, instead of ODR used by both A21 and B24, that allows for asymmetrical errors in the data and better accounting of outliers, which are highly prevalent in these datasets. For the Fe-MF comparison, A24 applied different statistical methods including a Bayesian model (), but it did not account for asymmetrical errors nor is the dataset weighted to account for outliers. B24 also applied OLS approach and obtained close to 1:1 relationship (), which in conjunction with their steep relation from ODR they interpreted as there not being enough evidence of a robust correlation. With a Bayesian approach, instead, we can inspect the posterior and identify any underlying discrepancies in the fit, thus quantifying the confidence in the relation as well as identify outliers, especially in light of a small data set. Therefore, we apply our nominal weighted fitting model (see Sec. II.4) to the A24 and B24 datasets for its added statistical advantage. We find that B24 dataset has a weaker slope to the one they report, which is closer to 1:1 , while the A24 dataset has a slope that is slightly shallower than their reported slope (see Fig. 6). This result indicates that both our and B24 datasets are fundamentally different compared to that of A24, which has a steeper slope. Furthermore, this is not attributed to the larger sample size of A24 (N=32) given the fact that when we limit our re-analysis of their data to the same exoplanets that are in our sample, we find the slope is still steep (). Thus, it is most likely due to the self-consistent stellar and planetary analysis done in this work.
III.4 Data needed to unequivocally test formation scenarios
@@
We use the GALAH dataset (Buder et al., 2024) to generate a population of stars and assume that each star hosts an exoplanet with some mass and corresponding radius such that the refractory ratios are exactly the same as the host star. Subsequently, we apply the current observational error scatter to this mock population (, for stars and exoplanets), and fit this Fe-MF star-planet mock data to a line to recover the slope and intercept. We repeat this process 1000 times to find the variance in these parameters. We find that the generated mock populations are fairly similar to our self-consistent exoplanet population. Our results for simulations with 21 planets in the sample (N) are shown in Figure 7. The most likely outcome is a population with no super-Mercuries – comprising half of the outcomes – although 1 or even 2 are still possible, accounting for 35 and 12 of the outcomes, respectively. We find instead that having or RTR-like planets is equally probable and irregardless of the number of outliers, we always recover a slope that is consistent with 1 and 0. Thus, the likelihood of randomly observing a population with 1 super-Mercury and 2-3 RTR-like exoplanets out of 21, which is what we have in our sample, is . The results from this mock population indicate that the larger range observed in the composition of exoplanets compared to that of stars may be explained, at the level, by the scatter produced due to observation uncertainty. It also means, however, it is entirely possible that the spread in planetary composition is real and reflects formation pathways away from a primordial origin. A larger sample and especially the prevalence or absence of super-Mercuries is the path to determine whether the spread in composition is real or an artifact of errors.
Next we examine the sample size needed to determine unequivocally a 1:1 relation given the current level of mass and radius certainty. Unsurprisingly but unfortunately, we find that a sample size of N, like the one we have, is unlikely to recover the true relationship () given the large uncertainty in the data – have m=0 within confidence (see Fig. 8). However, we find it is possible to robustly recover the correct best-fit line back, at 95 confidence when there are 150 planets. For a higher confidence level (3 ), we find that for N=150 there are around a quarter of runs with in the posterior. This means that to confidently constrain the composition using only Fe-MF of planets and stars, a sample that of at least 7 times the current size is needed. The good news is that PLATO is expected to deliver about 1000 planets with radii of after the first four mission years (Rauer et al., 2024).
@@
IV Discussion
IV.1 Super-Mercuries
We have shown the need to carefully select planetary masses and radius and in particular guarantee consistency in stellar parameters to ensure consistency in planetary M-R data. One of the shortcomings of using inconsistent data was the apparent existence of a group of super-Mercuries Plotnykov and Valencia (2020); Adibekyan et al. (2021); Schulze et al. (2021)) with implications on the existence of our own Mercury. In our solar system, Mercury holds a unique position as the innermost planet with an unusually large amount of iron, with Fe-MF Hauck et al. (2013) or 2-times enriched with respect to solar, making it an outlier compared to Earth, Venus and Mars which have similar refractory ratios to that of the Sun (Anders and Grevesse, 1989; Grevesse et al., 2007; Lodders et al., 2009; Asplund et al., 2009). There have been different proposed formation mechanisms (Ebel and Alexander, 2011; Hubbard, 2014; Cuello et al., 2016; Kruss and Wurm, 2020; Aguichine et al., 2020; Johansen and Dorn, 2022) to explain Mercury’s composition with giant impacts being the leading theory (Benz et al., 2007) albeit not without problems (Chau et al., 2018; Clement et al., 2021; Franco et al., 2022; Scora et al., 2024). It remains to be seen how unique its formation is. Thus, super-Mercuries are prime targets for understanding planetary formation and evolution theories.
In our updated dataset, the exoplanet M-R are revised such that we observe fewer super-Mercuries–only Kepler-107 c remains–and there is no evidence for a sub-population as seen by A21, in agreement with B24. This suggest the prevalence of fewer super-Mercuries in general, which we attribute to our homogeneous modeling and updates to the transit and RV observations. Nevertheless, we still find a few exoplanets–K2-38 b, K2-229 b and K2-291 b–that have relative iron-enrichment compared to their host star (Fe-MF). Some of these were previously thought super-Mercuries and would be important to follow up, given that there is no current formation theory to explain even this smaller iron-enrichment for massive rocky planets (Scora et al., 2020, 2024).
IV.2 Multi-Planet systems